### v1

@@ 1 Introduction
Contemporary LLMs possess the capability to comprehend natural language, discern user intent from input expressions, and execute tasks such as document summarization or translation [1], image generation [2] or code generation [3] tasks. Beyond these functions, LLMs also present the potential to deconstruct complex intents into discrete, actionable steps, thereby enabling the automated construction of workflows in a manner analogous to human reasoning [4]. LLMs have the potential to profoundly transform human-device interaction by supplanting rigid graphical interfaces with intuitive, conversational ones. Rather than navigating through menus or memorizing application-specific commands, users can articulate their objectives in natural language. LLMs are responsible for interpreting these inputs and orchestrating actions across various applications and services in a dynamic manner. As a consequence, complex tasks are simplified, and the system adapts to each user’s habits and context, thereby personalizing the user experience. This shift is not only particularly salient in the context of mobile devices, where screen space and input methods are constrained, but in other applications for human-computer interaction as well, such as robotics, where robots mimic human-like communication. Interfaces are undergoing a paradigm shift towards invisible, language-first systems, whereby interaction resembles conversing with a smart assistant more than utilising a conventional device.
@@
The prevailing design of operating systems has been predicated for the aforementioned interaction mechanisms with GUIs, hierarchical file management, and the shell, allocating the responsibility for interaction to the user. Therefore, the interaction mechanisms initiated by LLMs necessitate a reconceptualization of fundamental design decisions in contemporary operating systems. In our previous work, we presented the first step on the path to such a GUI-less operating system with the utilization of the proprietary gpt-4o-mini model [5]. Nevertheless, the proposed methodology engenders a considerable degree of interdependence. The advent of future mobile devices is poised to achieve user intentions independently of external infrastructure. It is imperative to incorporate LLMs into local devices to ensure autonomy, privacy, extensibility, and optimization. Open-source and open-access models hold considerable potential in this endeavor and can serve as pivotal elements, not only for the future integration of such LLMs on local devices, but for the development of future operating systems with open and transparent ecosystems. This leads to the research question that guides this study, which is as follows: "How effective are open-source and open-access models in resolving user intentions for future intent-based operating systems, and what areas of research and development are indicated to enable broad, multi-domain deployment?".
@@ 2 Methodologies and Approach
The collection of metrics is conducted in accordance with the experimental protocol, encompassing the Time to First Token, the Response Time, as well as the inclusion of preambles, postambles, and code comments for measuring, comparing, and objectively evaluating the model’s performance and responsiveness. The Time to First Token metric is defined as the duration required to receive the initial output from the specified LLM. This figure illustrates the model’s initialization and processing overhead prior to generation. On the other hand, the Response Time metric is defined as the total time required to receive the complete output, which should be measured in seconds and occur within a few seconds to avoid disrupting the user’s thought process [6]. Code elements, including comments, offer valuable insights into the decision-making process. However, these elements do so by increasing the response size, albeit to an infinitesimal degree. Preambles and postambles are integrated into the response and envelop the generated code block. These consist of explanations or introductory words from the LLM. The aforementioned elements are considered superfluous and serve only to augment the response size, thereby demonstrating an understanding of the designated role.
@@ 4 Discussion and Result Interpretation
Subsequent endeavors in this specific application domain pertain to the optimization of the aforementioned models for the purpose of further reducing the introduced system architecture. The efficacy of LLMs is contingent upon their incorporation into local devices. However, the substantial computational demands of the inference process currently necessitate execution on remote infrastructure. The processes of pruning, distillation, and quantization offer significant opportunities for the operation of LLMs at local scale. By decreasing the model size and computational demands without a substantial compromise in performance, these techniques enable the implementation of sophisticated AI models on devices with limited resources, not exclusive to mobile devices. Collectively, these technologies facilitate enhanced accessibility, reduced operational expenditures, augmented privacy measures, and expedited response times, unveiling novel prospects for real-world, on-device AI application and enabling the focus on operating system-oriented optimization of intent-based user interaction mechanisms. In this context, open-source and open-access models assume a particularly salient role. It is imperative to acknowledge that the reduction and optimization of the aforementioned models is not the sole pivotal step. While the employment of imperative programming languages as intermediate representations for workflows functions effectively in conjunction with LLMs, the necessity arises for an all-encompassing API to address the diverse user intents. This issue must be given due consideration for future research endeavors. Furthermore, the transition of authority and decision-making capacity to LLMs and AI in general gives rise to a substantial security concern, thereby prompting the exploration of critical research domains. The deliberate or inadvertent application of LLMs has the potential to result in adverse consequences. To illustrate this point, the direct execution of generated code in the system architecture under consideration introduces a security vulnerability. It is necessary to implement significant isolation and sandboxing mechanisms, as well as utilize operating system-provided capabilities. Another exemplary vector of attack in the experiments presented is targeted around the shell command, since it can be misused for direct access to the system, depending on the particular system configuration and setup. These findings align with the recent studies addressing the Shutdown Problem [7][8], which demonstrate that AI proactively undertakes measures to circumvent system shutdown. Alignment Faking [9] further illustrates how contemporary LLMs exhibit resistance to human intervention and correction. It is imperative to devise countermeasures and techniques to circumvent potential damage that may be engendered by the integration of LLMs and AI.
@@ 5 Related Work
In the seminal paper, the Transformer is introduced [10], which is a novel deep learning model founded on self-attention, which supersedes earlier Recurrent Neural Networks (RNNs) and Convolutional Neural Networks (CNNs). The Transformer model is distinguished by its parallelisability, accelerated training, and enhanced quality in Natural Language Processing [11][12]. This fundamental principle underlies the construction of all contemporary LLMs, including GPT, BERT, Falcon, Phi, and Qwen. A thorough examination of transformer design optimizations through the initial months of 2024 is elucidated in [13]. The overview encompasses FlashAttention-2, Mixture of Experts and Long Context Transformers. The widespread availability of ChatGPT [14] has led to a substantial increase in the number of applications under consideration. For instance, the application domains of public health and medicine have been the focus of study [15][16][17], as well as those of education and pedagogy [18][19]. The general applicability of AI necessitates its categorization, a subject that is addressed in [20]. Existing solutions such as Siri, Cortana as well as Alexa influenced the application domain of personal assistants. An overview of requirements of voice user interfaces, in particular for for blind and visually impaired users, is addressed by [21]. The integration of LLMs for machine-oriented user intention resolution is examined in [22] and [23], which also present AIOS, an operating system for LLM-based agents. [24] delineates a vision for AIOS as the core of future vehicle systems research. Recent research in this particular application domain includes the training of AI to directly operate existing GUI applications [25]. Subsequent research in the context of LLMs entails the investigation of the potential of LLMs to facilitate the recognition of user intentions within dialog systems [26]. A prototype tool that automatically generates business and scientific workflows using LLMs is presented in [27]. Another application of AI, particularly LLMs, is the generation of code, a subject that has been extensively studied. Tools such as GitHub Copilot provide assistance to engineers in routine tasks [28][29]. An evaluation of problem-solving through code generation with GPT language models has been conducted in [30] [31]. Common issues associated with the utilization of LLMs, including hallucinations and erroneous code generation, are addressed by techniques that are centered around fuzzing as well as static analysis [32] A number of novel approaches have been developed that utilize Grammar Augmentation [33] and the redesign of fundamental transformer decoding algorithms [3]. The utilization of AI and LLMs entails specific risks [34] and necessitates a systematic taxonomy of these risks, as outlined in [35]. Among the most critical aspects are privacy-related concerns associated with training data [36], as well as the handling of sensitive user information from communication platforms. A comprehensive survey on approaches to data privacy protection is provided in [37]. Further issues are related to linguistic biases [38].

### v2

@@ 1 Introduction
Contemporary LLMs possess the capability to comprehend natural language, discern user intent from input expressions, and execute tasks such as document summarization or translation [jin2024comprehensive], image generation [zhang2023controllable] or code generation [zhang2023planninglargelanguagemodels] tasks. Beyond these functions, LLMs also present the potential to deconstruct complex intents into discrete, actionable steps, thereby enabling the automated construction of workflows in a manner analogous to human reasoning [xie2024human]. LLMs have the potential to profoundly transform human-device interaction by supplanting rigid graphical interfaces with intuitive, conversational ones. Rather than navigating through menus or memorizing application-specific commands, users can articulate their objectives in natural language. LLMs are responsible for interpreting these inputs and orchestrating actions across various applications and services in a dynamic manner. As a consequence, complex tasks are simplified, and the system adapts to each user’s habits and context, thereby personalizing the user experience. This shift is not only particularly salient in the context of mobile devices, where screen space and input methods are constrained, but in other applications for human-computer interaction as well, such as robotics, where robots mimic human-like communication. Interfaces are undergoing a paradigm shift towards invisible, language-first systems, whereby interaction resembles conversing with a smart assistant more than utilising a conventional device.
@@
The prevailing design of operating systems has been predicated for the aforementioned interaction mechanisms with GUIs, hierarchical file management, and the shell, allocating the responsibility for interaction to the user. Therefore, the interaction mechanisms initiated by LLMs necessitate a reconceptualization of fundamental design decisions in contemporary operating systems. In our previous work, we presented the first step on the path to such a GUI-less operating system with the utilization of the proprietary gpt-4o-mini model [flerlage2025machinegeneratedcoderesolutionuser]. Nevertheless, the proposed methodology engenders a considerable degree of interdependence. The advent of future mobile devices is poised to achieve user intentions independently of external infrastructure. It is imperative to incorporate LLMs into local devices to ensure autonomy, privacy, extensibility, and optimization. Open-source and open-access models hold considerable potential in this endeavor and can serve as pivotal elements, not only for the future integration of such LLMs on local devices, but for the development of future operating systems with open and transparent ecosystems. This leads to the research question that guides this study, which is as follows: "How effective are open-source and open-access models in resolving user intentions for future intent-based operating systems, and what areas of research and development are indicated to enable broad, multi-domain deployment?".
@@ 2 Methodologies and Approach
The collection of metrics is conducted in accordance with the experimental protocol, encompassing the Time to First Token, the Response Time, as well as the inclusion of preambles, postambles, and code comments for measuring, comparing, and objectively evaluating the model’s performance and responsiveness. The Time to First Token metric is defined as the duration required to receive the initial output from the specified LLM. This figure illustrates the model’s initialization and processing overhead prior to generation. On the other hand, the Response Time metric is defined as the total time required to receive the complete output, which should be measured in seconds and occur within a few seconds to avoid disrupting the user’s thought process [nielsen1994usability]. Code elements, including comments, offer valuable insights into the decision-making process. However, these elements do so by increasing the response size, albeit to an infinitesimal degree. Preambles and postambles are integrated into the response and envelop the generated code block. These consist of explanations or introductory words from the LLM. The aforementioned elements are considered superfluous and serve only to augment the response size, thereby demonstrating an understanding of the designated role.
@@ 4 Discussion and Result Interpretation
Subsequent endeavors in this specific application domain pertain to the optimization of the aforementioned models for the purpose of further reducing the introduced system architecture. The efficacy of LLMs is contingent upon their incorporation into local devices. However, the substantial computational demands of the inference process currently necessitate execution on remote infrastructure. The processes of pruning, distillation, and quantization offer significant opportunities for the operation of LLMs at local scale. By decreasing the model size and computational demands without a substantial compromise in performance, these techniques enable the implementation of sophisticated AI models on devices with limited resources, not exclusive to mobile devices. Collectively, these technologies facilitate enhanced accessibility, reduced operational expenditures, augmented privacy measures, and expedited response times, unveiling novel prospects for real-world, on-device AI application and enabling the focus on operating system-oriented optimization of intent-based user interaction mechanisms. In this context, open-source and open-access models assume a particularly salient role. It is imperative to acknowledge that the reduction and optimization of the aforementioned models is not the sole pivotal step. While the employment of imperative programming languages as intermediate representations for workflows functions effectively in conjunction with LLMs, the necessity arises for an all-encompassing API to address the diverse user intents. This issue must be given due consideration for future research endeavors. Furthermore, the transition of authority and decision-making capacity to LLMs and AI in general gives rise to a substantial security concern, thereby prompting the exploration of critical research domains. The deliberate or inadvertent application of LLMs has the potential to result in adverse consequences. To illustrate this point, the direct execution of generated code in the system architecture under consideration introduces a security vulnerability. It is necessary to implement significant isolation and sandboxing mechanisms, as well as utilize operating system-provided capabilities. Another exemplary vector of attack in the experiments presented is targeted around the shell command, since it can be misused for direct access to the system, depending on the particular system configuration and setup. These findings align with the recent studies addressing the Shutdown Problem [thornley2024shutdown][thornley2024shutdown2], which demonstrate that AI proactively undertakes measures to circumvent system shutdown. Alignment Faking [greenblatt2024alignment] further illustrates how contemporary LLMs exhibit resistance to human intervention and correction. It is imperative to devise countermeasures and techniques to circumvent potential damage that may be engendered by the integration of LLMs and AI.
@@ 5 Related Work
In the seminal paper, the Transformer is introduced [vaswani2017attention], which is a novel deep learning model founded on self-attention, which supersedes earlier Recurrent Neural Networks (RNNs) and Convolutional Neural Networks (CNNs). The Transformer model is distinguished by its parallelisability, accelerated training, and enhanced quality in Natural Language Processing [kalla2023study][minaee2024large]. This fundamental principle underlies the construction of all contemporary LLMs, including GPT, BERT, Falcon, Phi, and Qwen. A thorough examination of transformer design optimizations through the initial months of 2024 is elucidated in [10.1145/3530811]. The overview encompasses FlashAttention-2, Mixture of Experts and Long Context Transformers. The widespread availability of ChatGPT [wubriefgpt] has led to a substantial increase in the number of applications under consideration. For instance, the application domains of public health and medicine have been the focus of study [biswas2023role][thirunavukarasu2023large][nazi2024large], as well as those of education and pedagogy [firat2023chat][kasneci2023chatgpt]. The general applicability of AI necessitates its categorization, a subject that is addressed in [morrisposition]. Existing solutions such as Siri, Cortana as well as Alexa influenced the application domain of personal assistants. An overview of requirements of voice user interfaces, in particular for for blind and visually impaired users, is addressed by [oumard2022pardon]. The integration of LLMs for machine-oriented user intention resolution is examined in [shi2024commands] and [mei2024aios], which also present AIOS, an operating system for LLM-based agents. [verhicle2024] delineates a vision for AIOS as the core of future vehicle systems research. Recent research in this particular application domain includes the training of AI to directly operate existing GUI applications [liu2024autoglmautonomousfoundationagents]. Subsequent research in the context of LLMs entails the investigation of the potential of LLMs to facilitate the recognition of user intentions within dialog systems [arora2024intentdetectionagellms]. A prototype tool that automatically generates business and scientific workflows using LLMs is presented in [xu2024llm4workflow]. Another application of AI, particularly LLMs, is the generation of code, a subject that has been extensively studied. Tools such as GitHub Copilot provide assistance to engineers in routine tasks [MORADIDAKHEL2023111734][wermelinger2023]. An evaluation of problem-solving through code generation with GPT language models has been conducted in [chen2021evaluating] [lin2024soen101codegenerationemulating]. Common issues associated with the utilization of LLMs, including hallucinations and erroneous code generation, are addressed by techniques that are centered around fuzzing as well as static analysis [Ouyang_2024] A number of novel approaches have been developed that utilize Grammar Augmentation [ugare2024syncodellmgenerationgrammar] and the redesign of fundamental transformer decoding algorithms [zhang2023planninglargelanguagemodels]. The utilization of AI and LLMs entails specific risks [bender2021dangers] and necessitates a systematic taxonomy of these risks, as outlined in [weidinger2022taxonomy]. Among the most critical aspects are privacy-related concerns associated with training data [carlini2021extracting], as well as the handling of sensitive user information from communication platforms. A comprehensive survey on approaches to data privacy protection is provided in [yan2024protecting]. Further issues are related to linguistic biases [fleisig2024linguistic].